OpenStudy (anonymous):

HELP ME ILL MEDAL Daniel watches 1 hour of television for every 3 hours of homework he does each week. If he did 12 hours of homework last week, how many hours of television did he watch? A. 12 B. 4 C. 36 D. 9

OpenStudy (aaronandyson):

Homework time TV Time 3 1 12 x

OpenStudy (anonymous):

its b ?

OpenStudy (aaronandyson):

3*x = 12*1 3x = 12 x = 12/3 x = ?

OpenStudy (aaronandyson):

Yes B

OpenStudy (aaronandyson):

12/3 = 4 so yes B :)